OpenStudy (anonymous):

The recommended dosage of the drug ampicillin for pediatric patients is 200 mg per kilogram of the patient's weight. How many mg of ampicillin should be given to a patient who weighs 39 pounds?

OpenStudy (anonymous):

1 lb = .454 kg \[\frac{200 mg ampicillin}{1 kg}\times\frac{.454 kg}{1 lb}\times\frac{39 lbs total body weight}{1 lb}\] Now cross-cancel all the units that you can and evaluate.

OpenStudy (anonymous):

Sorry, I forgot to use \textnormal

OpenStudy (texaschic101):

1 lb = .454 kg so 39 lbs would be (39)(.454) = 17.71 kg And 200 mg per kg = (200)(17.71) = your answer in mg I believe you can finish it from here :)